Do not use excel please, show work
Short Position Name: Date: Suppose that an investor sells 400 shares short at S50 per share. The initial margin is 50% and the maintenance margin is 30%. After 125 days, the investor purchases the shares for S40 and closes the short position. During the holding period, the shares paid a dividend of $2.50 per share. Show the calculations for the numbers that you enter in the T accounts a. Show the investors beginning T account (cales here) b. At what price would the investor receive a margin call? c. Show the investorsT account at the share price of $40. (cales, here d. What is the actual margin at a share price of $40? e. Calculate the investors return for the 125 day holding period
Answer #1

Value of investment= 400*50$= 20000$

Initial margin (50% of value)= 10000$

Minimum margin (30% of investment)= 6000$

a) Being the initial margin money transferred to broker account

Security Margin Account

Particulars Debit amount particulars Credit amount
To cash 10,000

Broker account

Particulars Debit amount particulars Credit amount

Cash Account/Bank account

Particulars Debit amount particulars Credit amount
by Security Margin 10000

b) Investor would receive a margin call when the price of the share increases by 10 $ or more i.e when the share price rise to 60$ or more he would receive a margin call.

c) Since the investor is short on shares and the price has fallen, he would have $ 4000 (10 $ *400 shares) added in his security margin account

Security Margin Account

Particulars Debit amount particulars Credit amount
To cash 10,000
to P/L A/c 4000

Profit & loss A/c

Particulars Debit amount particulars Credit amount
by margin 4000

d) Actual Margin is 14000$/400 shares= 35$

e) Investors return= 4000/10000*125/365= 13.70%

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW  

THANK YOU